the difference between two complementary angles is 14. what are the measures of both angles?
Answer by checkley77(12844)

X+X+14=90
2X=90-14
2X=76
X=76/2
X=38 ANS. FOR THE SMALLER ANGLE.
38+14=52 ANS. FOR THE LARGER ANGLE.
PROOF:
38+52=90
90=90